# Posterior Ramifications of Prior Dependence Structures

### 11 Dec 2023 v1

@@ 2.3 Illustration with Copula-Based Priors for the Multinomial Model
We now illustrate that prior dependence structures specified using the methods proposed by Elfadaly and Garthwaite (2017) and Wilson (2018) are routinely rejected when combined with the standard multinomial likelihood as data are observed. We consider the multinomial model with $w=3$ categories, which is parameterized by conditional probabilities $\boldsymbol{\theta}=(Z_{1},Z_{2})$. We specify $p(\boldsymbol{\theta})$ as in (4) where the marginal prior for $Z_{1}$ is $\text{BETA}(20,40)$, the marginal prior for $Z_{2}$ is $\text{BETA}(30,30)$, and these marginal priors are joined using a Gaussian copula parameterized with Pearson’s $\rho=-0.9$. This prior distribution conveys that we expect the multinomial probabilities to be roughly equal for all three categories. This prior specification could be facilitated using either of the methods by Elfadaly and Garthwaite (2017) and Wilson (2018). We note that the prior copula gives rise to a value for Kendall’s tau of $0.7132\text{sin}^{-1}(-0.9)/\pi=-$ 0.713.
@@
For illustration, we generate $10^{5}$ observations from the prior predictive distribution in (5) defined with $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ corresponding to the standard multinomial model and $p(\boldsymbol{\theta})$ as detailed in the previous paragraph. In Figure 2, we visualize the prior copula along with the posterior copula for $\boldsymbol{\theta}$ given $\boldsymbol{y}^{{}_{(n)}}$ corresponding to the first $n=10^{1},10^{2},10^{3}$, $10^{4}$, and $10^{5}$ observations from the generated sample. This figure shows that the posterior cannot retain the negative dependence structure present in the prior as data are observed from the prior predictive distribution. Even though the prior $p(\boldsymbol{\theta})$ accurately characterizes the data generation process for the multinomial data, the multinomial likelihood function cannot accommodate this dependence between $Z_{1}$ and $Z_{2}$. As the sample size increases, we observe that the impact of the likelihood function on the posterior overwhelms that of the prior. We may require a very large multinomial sample, but the prior dependence structure will eventually be rejected when the conditions for Theorem 1 hold.
Figure 2: Rejection of negative prior dependence in the joint posterior of $Z_{1}$ and $Z_{2}$ at sample sizes $n=10^{1},10^{2},10^{3}$, $10^{4}$, and $10^{5}$.
@@
The copula-based priors proposed by Elfadaly and Garthwaite (2017) and Wilson (2018) define valid posteriors for $\boldsymbol{\theta}$ when combined with the multinomial likelihood, but Figure 2 corroborates that we would draw vastly different conclusions about the posterior dependence structure of $\boldsymbol{\theta}$ for small and large samples. Elfadaly and Garthwaite (2017) provided R code to combine their priors with the multinomial likelihood function, and Wilson et al. (2021) stated that these priors could be used as prior distributions in Bayesian analyses. However, neither contribution clearly disclosed that these more flexible prior dependence structures cannot be retained as multinomial data are collected. Perhaps the posterior should be defined using a likelihood function for categorical data that allows for this dependence if these more flexible dependence structures are plausible a priori. This work does not consider how such modifications might allow for more flexible dependence structures to be retained a posteriori. Instead, we alert practitioners to situations where chronic rejection of prior dependence occurs. The adequacy stage of the prior elicitation framework proposed by Garthwaite et al. (2005) involves examining the prior’s fitness for purpose. Given a candidate prior dependence structure, Corollary 2 helps assess whether $p(\boldsymbol{\theta})$ is a suitable analysis prior. We note that prior distributions may be elicited to serve different purposes, and we elaborate on these situations in Section 3.
3 Design Priors
@@
We clarify that the flexible dependence structures facilitated by copula-based priors are useful in certain situations. Not all prior distributions are elicited with the intent of defining a posterior. Prior distributions are regularly used for design purposes. For instance, priors might be used to summarize expert opinion to choose inputs for a decision model (Garthwaite et al., 2005). Design priors are also used in experimental settings to conduct sample size determination (De Santis, 2007; Berry et al., 2011; Gubbiotti and De Santis, 2011). In these settings, design priors are often informative and concentrated on $\boldsymbol{\theta}$ values that are relevant to the objective of the study. The prior $p_{D}(\boldsymbol{\theta})$ from (5) could be considered as a design prior in the context of this paper. The design prior is sometimes called the sampling prior, whereas the analysis prior $p(\theta)$ in (4) is sometimes named the fitting prior (Wang and Gelfand, 2002; Sahu and Smith, 2006).
Wilson (2018) illustrated the elicitation method with vine copulas using a multinomial model that characterized various states for the capacity of a complex engineering system. It is possible that the additional flexibility in the copula-based prior allows for subject matter expertise to be more accurately characterized. Perhaps this improved accuracy could give rise to better mitigation of safety implications associated with the engineering system. In that case, there may be value in taking the extra effort to specify a vine copula. If this more flexible prior were not combined with the multinomial likelihood function, there would be no concerns with such an approach. However, it is difficult for practitioners to make informed decisions about whether specifying a prior copula is worthwhile when it is not transparent that this additional effort could undermine the credibility of their posterior analysis. Copula-based prior elicitation procedures – such as those proposed by Elfadaly and Garthwaite (2017) and Wilson (2018) – are therefore useful in design settings, but conditions for their responsible use should be more clearly defined.
@@ 4.1 Convergence of the Posterior Mode
For an arbitrary model, it may be challenging to correctly specify the dependence structure a priori. As such, we consider how the choice of copula for the prior distribution impacts the posterior. We suppose that two potential priors for $\boldsymbol{\theta}$, denoted $p_{1}(\boldsymbol{\theta})$ and $p_{2}(\boldsymbol{\theta})$, are defined as in (4) using the same marginal distributions $F_{1},...,F_{d}$ but different copula density functions $c_{1}(\boldsymbol{u})$ and $c_{2}(\boldsymbol{u})$. We require that both copula density functions are absolutely continuous and twice differentiable with respect to $\boldsymbol{u}=(u_{1},...,u_{d})$. To ensure the domain of the parameter space is not inadvertently restricted, the priors should be chosen such that $p_{1}(\boldsymbol{\theta})>0$ and $p_{2}(\boldsymbol{\theta})>0$ for all $\boldsymbol{\theta}\in\boldsymbol{\Theta}$.
We define posteriors $p_{1}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ and $p_{2}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ by combining the likelihood $L(\boldsymbol{\theta};\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ for the model $m(\boldsymbol{y}^{{}_{(n)}}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ with $p_{1}(\boldsymbol{\theta})$ and $p_{2}(\boldsymbol{\theta})$, respectively. We summarize each posterior via its posterior mode for $\boldsymbol{\theta}$, denoted by $\tilde{\boldsymbol{\theta}}^{{}_{(k)}}=\text{arg\,max}_{{\boldsymbol{\theta}}}~{}p_{k}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ for $k=1,2$. In this section, we consider the convergence of the posterior mode to a fixed value $\boldsymbol{\theta}_{0}$. Unlike in Theorem 1, we exclusively consider degenerate priors $p_{D}(\boldsymbol{\theta})$ for the prior predictive distribution of $\boldsymbol{Y}^{{}_{(n)}}$ in (5). This restriction is necessary for the posterior of $\boldsymbol{\theta}$ to converge around the fixed value $\boldsymbol{\theta}_{0}$. This degenerate prior $p_{D}(\boldsymbol{\theta})$ is distinct from the analysis priors $p_{1}(\boldsymbol{\theta})$ and $p_{2}(\boldsymbol{\theta})$ used to define the relevant posteriors. For a given sample $\boldsymbol{y}^{{}_{(n)}}$, we compare the Euclidean distance between $\boldsymbol{\theta}_{0}$ and each posterior mode, denoted by $\mathcal{D}_{k}=\lVert\tilde{\boldsymbol{\theta}}^{{}_{(k)}}-\boldsymbol{\theta}_{0}\rVert_{2}$ for $k=1,2$. Theorem 2 indicates that we generally do not expect the choice of prior copula to impact whether the posterior mode is closer to $\boldsymbol{\theta}_{0}$ for large sample sizes $n$.
@@
Theorem 2 explains that whether the choice of prior copula gives rise to faster convergence depends on the function $\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))$. While $c_{1}(\cdot)$ and $c_{2}(\cdot)$ are functions of $\boldsymbol{\theta}$ because $\boldsymbol{u}=(F_{1}(\theta_{1}),...,F_{d}(\theta_{d}))$, we consider the partial derivatives of the copula log-density functions with respect to $\boldsymbol{u}$ instead of $\boldsymbol{\theta}$. For $j=1,...,d$, it follows by the chain rule that
@@
$\begin{split}\frac{\partial}{\partial\theta_{j}}\bigg{[}\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))\bigg{]}=\\ f_{j}(\theta_{j})\frac{\partial}{\partial u_{j}}\bigg{[}\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))\bigg{]}.\end{split}$ (7)

### 23 Mar 2024 v2

@@ 2.3 Illustration with Copula-Based Priors for the Multinomial Model
We now apply Corollary 1 with the multinomial model. In Appendix A of the supplementary materials, we show that the inverse Fisher information matrix for the multinomial model parameterized in terms of $\boldsymbol{\theta}=(Z_{1},Z_{2},\dots,Z_{w-1})$ from (2) is diagonal for all possible $(Z_{1},Z_{2},\dots,Z_{w-1})\in(0,1)^{w-1}$. If $p(\boldsymbol{\theta})$ incorporates any positive or negative dependence between the conditional multinomial probabilities, this dependence structure is therefore a chronically rejected one.
@@
For $w=3$ categories, we illustrate this phenomenon when combining such priors with the standard multinomial likelihood, which is parameterized by conditional probabilities $\boldsymbol{\theta}=(Z_{1},Z_{2})$. We specify $p(\boldsymbol{\theta})$ as in (4) where the marginal prior for $Z_{1}$ is $\text{BETA}(20,40)$, the marginal prior for $Z_{2}$ is $\text{BETA}(30,30)$, and these marginal priors are joined using a Gaussian copula parameterized with Pearson’s $\rho=-0.9$. This prior distribution conveys that we expect the multinomial probabilities to be roughly equal for all three categories. This prior specification could be facilitated using either of the methods by Elfadaly and Garthwaite (2017) or Wilson (2018). We note that the prior copula gives rise to a value for Kendall’s $\tau$ of $0.7132\text{sin}^{-1}(-0.9)/\pi=-$ 0.713.
For each of 10000 simulation repetitions, we generated $\boldsymbol{\theta}_{0}=(Z_{1,0},Z_{2,0})$ from the prior specified in the previous paragraph. We generated $\boldsymbol{Y}^{{}_{(n)}}$ for $n=10$ from the multinomial model parameterized by $\boldsymbol{\theta}_{0}$. For each sample, we approximated the posterior of $\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}$ using sampling-resampling methods (Rubin, 1988), where the proposal distribution was the posterior of $\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}$ obtained by independently joining the marginal priors for $Z_{1}$ and $Z_{2}$. For each posterior sample, we estimated Kendall’s $\tau$ for $Z_{1}$ and $Z_{2}$. This process was repeated for $n=\{10^{2},10^{3},10^{4},10^{5}\}$. The range of Kendall’s $\tau$ values observed in this numerical study is summarized in Table 1.
Table 1: Kendall’s $\tau$ values for 10000 posteriors of $Z_{1}$ and $Z_{2}$ across various sample sizes $n$
Kendall’s $\tau$
$n$ Minimum Median Maximum
$10^{1}$0.7128 -0.6806 -0.6638
$10^{2}$0.5488 -0.5049 -0.4815
$10^{3}$0.2103 -0.1611 -0.1288
$10^{4}$0.0464 -0.0214 0.0062
$10^{5}$0.0326 -0.0022 0.0247
@@
This table shows that the posterior cannot retain the negative dependence structure present in the prior as data are observed from the prior predictive distribution. Even though the analysis prior $p(\boldsymbol{\theta})$ coincides with that used to define the prior predictive distribution of $\boldsymbol{Y}^{{}_{(n)}}$, the multinomial likelihood function cannot accommodate this dependence between $Z_{1}$ and $Z_{2}$. As the sample size increases, we observe that the impact of the likelihood function on the posterior overwhelms that of the prior, and the prior dependence structure will eventually not be retained when the conditions for Corollary 1 hold.
@@
The copula-based priors proposed by Elfadaly and Garthwaite (2017) and Wilson (2018) define valid posteriors for $\boldsymbol{\theta}$ when combined with the multinomial likelihood, but Table 1 corroborates that we would draw vastly different conclusions about the posterior dependence structure of $\boldsymbol{\theta}$ for small and large samples. Elfadaly and Garthwaite (2017) provided R code to combine their priors with the multinomial likelihood function, and Wilson et al. (2021) stated that these priors could be used for Bayesian analysis. However, neither contribution clearly disclosed that these more flexible prior dependence structures cannot be retained as multinomial data are collected. Given a candidate prior dependence structure, Corollary 1 helps assess whether this dependence structure for $\boldsymbol{\theta}$ can be retained a posteriori. We elaborate on how to determine whether the inability to retain the prior dependence structure presents practical issues in Section 3.
3 Objectives for Dependence Structure Specification
This section outlines several objectives that practitioners may wish to achieve when specifying prior distributions and their dependence structures. For each objective, we discuss whether retention of the prior dependence structure as defined in Corollary 1 is important. This knowledge can be used to assess the adequacy of an elicited prior distribution.
3.1 Supplementation of Small Samples
Even if a prior dependence structure cannot be retained as the sample size $n$ increases, it may be used to supplement information from small samples. The following example that illustrates the utility of chronically rejected dependence structures for small samples was helpfully provided by a reviewer of this paper. We suppose that a practitioner aims to model the heights of a new animal species with a $\mathcal{N}(\mu,\sigma^{2})$ distribution. The practitioner does not know the typical heights for members of this species, but it is reasonable to expect that $\mu$ and $\sigma$ are on the same scale. Hence, positive prior dependence between $\mu$ and $\sigma^{2}$ is sensible. Upon observing a single member of this species, the practitioner can estimate $\mu$ – even though the single observation does not directly provide information about $\sigma^{2}$. If $\mu$ and $\sigma^{2}$ were positively correlated a priori, the posterior could convey that a value of one nanometer for $\sigma$ would be unlikely given an observed height of one meter.
For this model with $\boldsymbol{\theta}=(\mu,\sigma^{2})$, the inverse Fisher information matrix $\mathcal{I}(\boldsymbol{\theta})^{-1}$ is diagonal for all possible values of $\mu\in\mathbb{R}$ and $\sigma^{2}\in\mathbb{R}^{+}$. It follows by Corollary 1 that the joint posterior of $\mu$ and $\sigma^{2}$ cannot retain a prior dependence structure for which Kendall’s $\tau$ is not 0. Under the conditions for Corollary 1, any attempt to inject positive or negative dependence into this posterior will be unsuccessful in the limit of infinite data. That is, once both $\mu$ and $\sigma^{2}$ are precisely estimated from the data, there will be no practical correlation between their estimates. However, imposing prior dependence is nevertheless useful for small sample sizes.
While chronically rejected dependence structures can supplement information from small samples, it would be ideal if the source of posterior dependence for $\boldsymbol{\theta}$ were transparent. Corollary 1 can be used to determine whether the prior or likelihood function gives rise to this dependence. This knowledge would provide additional context with which to interpret posterior analyses. This example also suggests that it is not worthwhile to spend time eliciting complicated prior dependence structures that cannot be retained when collecting large samples.
The use of informative priors to improve model identifiability is related to supplementing small samples with prior information. Although unidentifiability arises in other contexts, the sample size is often small with respect to the complexity of models that cannot be fully identified. Previous contributions detailed how using informative priors to enhance model identifiability impacts model interpretation and the convergence of Markov chain Monte Carlo methods (Gelfand and Sahu, 1999; Eberly and Carlin, 2000; Gustafson, 2005). These contributions mainly considered informative marginal priors, and discussion of informative prior dependence structures was limited.
3.2 Coverage of Credible Sets
Prior dependence structures may also be specified to calibrate Bayesian credible sets. In this context, Gustafson (2012) referred to the data generating prior as “nature’s” prior. The data generation prior in our framework is $p_{D}(\boldsymbol{\theta})$ from (5). If the analysis prior $p(\boldsymbol{\theta})$ coincides with “nature’s” prior and the model $m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ is correctly specified, credible sets attain their nominal coverage (Gustafson, 2012). We let $\mathcal{J}_{\boldsymbol{\theta},1-\alpha}(\boldsymbol{y}^{{}_{(n)}})$ denote a credible set for $\boldsymbol{\theta}$ given the observed data $\boldsymbol{y}^{{}_{(n)}}$ with coverage $1-\alpha$. A popular choice for the credible set is that of highest posterior density (HPD). Credible sets attain their nominal coverage when
$Pr(\boldsymbol{\theta_{0}}\in\mathcal{J}_{\boldsymbol{\theta},1-\alpha}(\boldsymbol{Y}^{{}_{(n)}}))=1-\alpha,$ (6)
where $\boldsymbol{Y}^{{}_{(n)}}\sim m(\hskip 1.0pt\cdot\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta}_{0})$ such that $\boldsymbol{\theta}_{0}$ is drawn from $p_{D}(\boldsymbol{\theta})$. The probabilistic statement in (6) is therefore made with respect to repeated sampling from the prior predictive distribution of $\boldsymbol{Y}^{{}_{(n)}}$. Once $\boldsymbol{y}^{{}_{(n)}}$ is observed, the credible set is not random and either contains or does not contain a given value $\boldsymbol{\theta_{0}}$.
We now investigate how the prior dependence structure impacts the calibration of credible sets for the multinomial example from Section 2.3. We used the same prior predictive distribution of $\boldsymbol{Y}^{{}_{(n)}}$ for this numerical study, joining the $\text{BETA}(20,40)$ prior for $Z_{1}$ and $\text{BETA}(30,30)$ prior for $Z_{2}$ with a Gaussian copula parameterized with Pearson’s $\rho=-0.9$. For each of 10000 simulation repetitions, we approximated the posterior of $\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}}$ as described in Section 2.3 with “nature’s” analysis prior. For each posterior, we approximated its 95% HPD set for $Z_{1}$ and $Z_{2}$ using two-dimensional kernel density estimation (Ripley, 2002). Empirical coverage was estimated as the proportion of simulation repetitions for which the parameter value $\boldsymbol{\theta}_{0}=(Z_{1,0},Z_{2,0})$ used to generate the multinomial data was contained in this HPD set. We implemented this process for $n=\{10^{1},10^{2},10^{3},10^{4},10^{5}\}$. We then repeated this process for analysis priors $p(\boldsymbol{\theta})$ that joined the marginal beta priors from “nature’s” prior with a Gaussian copula parameterized by Pearson’s $\rho=\{-0.95,-0.85,-0.80,\dots,0.95\}$. The results from this numerical study are visualized in Figure 3.
Figure 3: Empirical coverage of 95% HPD sets for the multinomial parameter $\boldsymbol{\theta}=(Z_{1},Z_{2})$ across 10000 posteriors. The horizontal dotted line denotes the nominal coverage, and the vertical one denotes “nature’s” prior.
The fact that all five colored curves in Figure 3 roughly intersect the horizontal dotted line when $\rho=-0.9$ confirms the result in (6). For small sample sizes $n$, empirical coverage generally deviated from the nominal coverage when $\rho\neq-0.9$. These deviations resulted in worse coverage when $p(\boldsymbol{\theta})$ was specified such that the magnitude of the negative dependence between $Z_{1}$ and $Z_{2}$ was overstated or when the direction of the dependence was inaccurate. The empirical coverage was greater than the nominal coverage when $p(\boldsymbol{\theta})$ incorporated slightly weaker negative dependence between $Z_{1}$ and $Z_{2}$ than “nature’s” prior. For those settings, the copula density function for $p(\boldsymbol{\theta})$ is flatter on $[0,1]^{2}$ than the copula density function that defines “nature’s” prior. The resulting credible sets cover a greater range of $\boldsymbol{\theta}$ values but still account for the correct direction of dependence between $Z_{1}$ and $Z_{2}$.
As the sample size $n$ increases, the impact of prior dependence is reduced and the empirical coverage approaches the nominal coverage for all $\rho$ values considered. This conclusion is to be expected given the asymptotic theory discussed in Section 2. The insights drawn from Figure 3 are not solely applicable to chronically rejected prior dependence structures. We observed similar results for a numerical study with the gamma model involving analysis priors that do not satisfy the conditions for chronic rejection in Corollary 1. These results are detailed in Appendix B of the online supplement.
Given the results from this subsection, we suggest erring on the side of understating the strength of dependence structures a priori. This recommendation is not made to encourage practitioners to engineer credible sets with coverages that exceed their nominal values. This excess coverage may not be desirable and cannot persist as the sample size increases. We instead make this recommendation to mitigate the potential ramifications of overstating the strength of dependence structures on the calibration of credible sets. The notion of harmful priors has been discussed elsewhere (see e.g., Reimherr et al. (2021)). However, priors that are harmful to the calibration of credible sets are not necessarily harmful priors in alternative contexts.
3.3 Inference Regarding Dependence Structures
One may also conduct posterior analyses to draw conclusions about the dependence structure of $\boldsymbol{\theta}$. The dependence structure is also relevant when considering the posterior of a function of several parameters in $\boldsymbol{\theta}$. To achieve these objectives with posterior analysis, we recommend using Corollary 1 to determine whether the prior dependence structure can be retained. Per Table 1, we can draw vastly different conclusions about chronically rejected dependence structures given small and large samples obtained from the same data generation process. Corollary 1 provides context to help gauge the reliability of our conclusions for small samples. For large samples, Corollary 1 dictates if time should be invested eliciting a dependence structure based on whether it can be retained a posteriori.
Corollary 1 could also be used to determine whether it is sensible to make inferences about dependence structures. While copula-based priors may be able to accommodate flexible dependence structures, the statistical models chosen for the likelihood function may not have this capability. For the multinomial model, there is a single component for each observation of $\boldsymbol{y}^{{}_{(n)}}=\{y_{i}\}_{i=1}^{n}$. It would be difficult to specify a likelihood function that could accommodate complex dependence structures between the conditional multinomial probabilities given the available data. As such, Corollary 1 could diagnose philosophical issues with objectives for posterior analyses.
If each observation of $\boldsymbol{y}^{{}_{(n)}}$ is comprised of multiple components, incorporating a copula into the likelihood function may promote greater coherence between the prior and posterior dependence structures. For illustration, we suppose that $\boldsymbol{y}^{{}_{(n)}}=\{(y_{i},y^{*}_{i})\}_{i=1}^{n}$, where $y_{i}\sim\text{EXP}(\lambda)$, $y^{*}_{i}\sim\text{EXP}(\kappa)$, and $\lambda$ and $\kappa$ are rates. The likelihood function for this example is such that
$\begin{split}L(\boldsymbol{\theta};\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})\propto\prod_{i=1}^{n}&\lambda e^{-\lambda y_{i}}\times\kappa e^{-\kappa y^{*}_{i}}\times\\ &c(1-e^{-\lambda y_{i}},1-e^{-\kappa y^{*}_{i}};\boldsymbol{\upsilon}),\end{split}$ (7)
where the copula density function is parameterized by $\boldsymbol{\upsilon}$ and $\boldsymbol{\theta}=(\lambda,\kappa,\boldsymbol{\upsilon})$. If $c(u_{1},u_{2};\boldsymbol{\upsilon})$ corresponds to the independence copula, the posterior correlation between $\lambda$ and $\kappa$ based on the inverse Fisher information will approach 0 as $n\rightarrow\infty$. More flexible dependence structures for $\lambda$ and $\kappa$ could be accommodated a posteriori given different choices for $c(u_{1},u_{2};\boldsymbol{\upsilon})$. In those settings, eliciting dependence structures for $\lambda$ and $\kappa$ could be worthwhile – even for large samples since the dependence structure might be retained.
3.4 Design Priors
@@
Not all prior distributions are elicited with the intent of defining a posterior. Prior distributions are regularly used for design purposes. For instance, priors might be used to summarize expert opinion to choose inputs for a decision model (Garthwaite et al., 2005). Design priors are also used in experimental settings to conduct sample size determination (De Santis, 2007; Berry et al., 2011; Gubbiotti and De Santis, 2011). In these settings, design priors are often informative and concentrated on $\boldsymbol{\theta}$ values that are relevant to the objective of the study. Data generated according to the design prior are often combined with an uninformative analysis prior to assess whether a posterior criterion is satisfied. The prior $p_{D}(\boldsymbol{\theta})$ from (5) could be considered as a design prior in the context of this paper.
Design priors are not directly combined with a likelihood function. It is therefore not an issue if the dependence structure in the design prior satisfies the conditions for chronic rejection outlined in Corollary 1. Generally, we do not need to make separate considerations for small and large samples when using design priors. It is possible that the design prior might coincide with an analysis prior $p(\boldsymbol{\theta})$ that defines a posterior of $\boldsymbol{\theta}$. In that event, the design prior is subject to the previous recommendations in this section. The discussion on the coverage of credible sets in Section 3.2 may be relevant if the design and analysis priors do not coincide.
3.5 Posterior Concentration
Recent work has suggested that the choice of prior dependence structure can expedite the convergence of the posterior around a fixed parameter value $\boldsymbol{\theta}_{0}$. In particular, Michimae and Emura (2022) suggested that joint priors based on Archimedean copulas lead to more accurate and concentrated posterior distributions in the context of Bayesian ridge regression. The accuracy and concentration of the posterior of $\boldsymbol{\theta}$ were considered via the total mean absolute error between the posterior median and a fixed parameter value $\boldsymbol{\theta}_{0}$. Michimae and Emura’s (2022) numerical studies showed that the posterior was more concentrated around $\boldsymbol{\theta}_{0}$ when the marginal priors for the regression coefficients were joined using (Archimedean) Clayton and Gumbel copulas (Nelsen, 2006) instead of more standard Gaussian copulas.
The asymptotic theory presented in Section 2 informally suggests the choice of prior dependence structure cannot give rise to increased posterior concentration for large samples. For small samples, further investigation into how the prior dependence structure prompts increased posterior concentration is required. This investigation would disclose whether improving posterior concentration is a sensible objective for prior dependence specification. These recent recommendations serve as one source of motivation to study the impact of the prior copula on the posterior distribution, which we do in Section 4.
@@ 4.1 Convergence of the Posterior Mode
For an arbitrary model, it may be challenging to correctly specify the dependence structure a priori. As such, we consider how the choice of copula for the prior distribution impacts the posterior. We suppose that two potential priors for $\boldsymbol{\theta}$, denoted $p_{1}(\boldsymbol{\theta})$ and $p_{2}(\boldsymbol{\theta})$, are defined as in (4) using the same marginal distributions $F_{1},\dots,F_{d}$ but different copula density functions $c_{1}(\boldsymbol{u})$ and $c_{2}(\boldsymbol{u})$. We require that both copula density functions are absolutely continuous and twice differentiable with respect to $\boldsymbol{u}=(u_{1},\dots,u_{d})$. To ensure the domain of the parameter space is not inadvertently restricted, the priors should be chosen such that $p_{1}(\boldsymbol{\theta})>0$ and $p_{2}(\boldsymbol{\theta})>0$ for all $\boldsymbol{\theta}\in\boldsymbol{\Theta}$.
We define posteriors $p_{1}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ and $p_{2}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ by combining the likelihood $L(\boldsymbol{\theta};\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ for the model $m(\boldsymbol{y}^{{}_{(n)}}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{\theta})$ with $p_{1}(\boldsymbol{\theta})$ and $p_{2}(\boldsymbol{\theta})$, respectively. We summarize each posterior via its posterior mode for $\boldsymbol{\theta}$, denoted by $\tilde{\boldsymbol{\theta}}^{{}_{(k)}}=\text{arg\,max}_{{\boldsymbol{\theta}}}~{}p_{k}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ for $k=1,2$. In this section, we consider the convergence of the posterior mode to a fixed value $\boldsymbol{\theta}_{0}$. For a given sample $\boldsymbol{y}^{{}_{(n)}}$, we compare the Euclidean distance between $\boldsymbol{\theta}_{0}$ and each posterior mode, denoted by $\mathcal{D}_{k}=\lVert\tilde{\boldsymbol{\theta}}^{{}_{(k)}}-\boldsymbol{\theta}_{0}\rVert_{2}$ for $k=1,2$. Theorem 2 indicates that we generally do not expect the choice of prior copula to impact whether the posterior mode is closer to $\boldsymbol{\theta}_{0}$ for large sample sizes $n$.
@@
Theorem 2 explains that whether the choice of prior copula gives rise to faster convergence depends on the function $\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))$. The posterior mode $\tilde{\boldsymbol{\theta}}^{{}_{(k)}}$ maximizes the logarithm of $p_{k}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})$ for $k=1,2$. The log-posteriors $\text{log}\left[p_{1}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})\right]$ and $\text{log}\left[p_{2}(\boldsymbol{\theta}\hskip 1.0pt|\hskip 1.0pt\boldsymbol{y}^{{}_{(n)}})\right]$ differ only by their prior copula log-density functions. Differences in $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$ and $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$ are therefore driven by differences in $\log(c_{1}(\boldsymbol{u}))$ and $\log(c_{2}(\boldsymbol{u}))$. Each copula log-density function prompts an additive contribution to the log-posterior. We suppose that $\boldsymbol{u}_{0}$ is not a stationary point of $\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))$ in part $(a)$. As $n$ increases, the contribution from $\log(c_{2}(\boldsymbol{u}))$ will not uniformly force its posterior mode closer to (or further from) $\boldsymbol{\theta}_{0}$ for all samples $\boldsymbol{y}^{{}_{(n)}}$ than that from $\log(c_{1}(\boldsymbol{u}))$. When $\boldsymbol{u}_{0}$ is instead a local maximum of $\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))$, the contribution from $\log(c_{2}(\boldsymbol{u}))$ will force $\tilde{\boldsymbol{\theta}}^{{}_{(2)}}$ closer to $\boldsymbol{\theta}_{0}$ than $\tilde{\boldsymbol{\theta}}^{{}_{(1)}}$ for all samples $\boldsymbol{y}^{{}_{(n)}}$ as $n$ increases. This case is considered in part $(b)$.
@@
While $c_{1}(\cdot)$ and $c_{2}(\cdot)$ are functions of $\boldsymbol{\theta}$ because $\boldsymbol{u}=(F_{1}(\theta_{1}),\dots,F_{d}(\theta_{d}))$, we consider the partial derivatives of the copula log-density functions with respect to $\boldsymbol{u}$ instead of $\boldsymbol{\theta}$. For $j=1,\dots,d$, it follows by the chain rule that
@@
$\begin{split}\frac{\partial}{\partial\theta_{j}}\bigg{[}\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))\bigg{]}=f_{j}(\theta_{j})\frac{\partial}{\partial u_{j}}\bigg{[}\text{log}(c_{2}(\boldsymbol{u}))-\text{log}(c_{1}(\boldsymbol{u}))\bigg{]}.\end{split}$ (8)